[Brahms-l] Shift report 20050323 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Mar 23 2005 - 07:59:50 EST
Shifter leader was SB
Shifters were 


1. Five stores, each about an hour and a half long
------------------------------------------------------
2. Run No. --   Fs @8 deg,@B430 -- MRS@40 deg, @B700
------------------------------------------------------
   13942       trig2(3.7k)           trig3(65k)
-------------------------------------------------------
   Run No. --   Fs @4 deg,@B861 -- MRS@40 deg, @B700
-------------------------------------------------------
   13943       trig2(20k)           trig3(36k)
   13944       trig2(8.8k)          trig3(34k)
   13945       trig2(6.1k)          trig3(26k)
   13946       trig2(21k)           trig3(73k)
   13947       trig2(13k)           trig3(50k)
   13948       trig2(19k)           trig3(67k)
   13949       trig2(13k)           trig3(53k)
   13950       trig2(23k)           trig3(83k)
   13951       trig2(10k)           trig3(43k)
   13952       trig2(27k)           trig3(49k)
   13953       running 
-------------------------------------------------
3. No detector problems so far


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 23 08:00:08 2005

This archive was generated by hypermail 2.1.8 : Wed Mar 23 2005 - 08:00:20 EST